# Dense Matter in Neutron Stars with eXTP

### v1

@@ 2 Pulse profile modeling
Pulse profile modeling (PPM) exploits the variation in brightness and spectrum as a function of the rotational phase to extract properties of either the emission or the NS itself. In the context of constraints on dense matter, PPM is used to model the emission of MSPs (rotation- or accretion- powered) [272]: more specifically the emission originating close to the NS surface, i.e., the emission affected the most by the strong NS surface gravity. As this is related to the NS compactness, PPM facilitates the measurement of the NS mass and radius, among other properties. This requires not only high quality pulse profile data (e.g., with order 10–100 sec timing resolution and 100 eV spectral resolution in the soft X-ray band) but also the ability to simulate pulse profiles using an adequate physical model (to form the basis for likelihood evaluations). Model ingredients include the physics of relativistic ray-tracing in the space-time of a rapidly rotating NS for a given mass and radius, a prescription for the possible properties of the X-ray emitting hotspots, an atmospheric beaming function, the effects of interstellar absorption and the instrument response (taking into account calibration uncertainty). Other relevant parameters will include the distance, pulsar inclination and any non-source background. Priors must be chosen for all of the model parameters. For a detailed overview of the simulation and inference process involved in pulse profile modeling, see [273, 274] and references therein.
@@ 2.1.1 State of the art - MSP PPM with NICER
NICER is a soft X-ray telescope operating on the International Space Station since 2017. Its primary science goal is to explore the EOS of NSs by collecting high quality pulse profile data sets of MSPs to be used for PPM. The NICER MSPs are also radio pulsars: radio timing provides the spin ephemeris necessary to generate the X-ray pulse profile, and other informative priors such as distance or, if the source is in a binary, the mass and inclination [282, 283]. Constraints on background can come either from NICER background models (e.g., [284]) or indirectly via joint fitting of (phase-averaged) data for the same source from other X-ray observatories with imaging capabilities (e.g., XMM-Newton, with a well constrained background, enabling a good estimate of the number of photons coming from the source in the XMM-Newton data set). To date, PPM using NICER data has yielded results for four MSPs: PSR J0030+0451, for which the mass and inclination are not known a priori [285, 286, 277, 287, 288, 289, 290]; the 2.1 pulsar PSR J0740+6620 [291, 292, 293, 294, 287, 295, 296, 297]; the 1.4 pulsar PSR J04374715 [281]; and PSR J12311411, which has only a weakly-informative prior on mass and inclination [298, 275]. In support of this effort, extensive work has gone into developing and testing PPM pipelines [273, 274, 299, 300]. The tightest constraints on radius so far (for PSR J04374715, [281]) are at the % level (68% credible interval). The results are being used to place constraints on both the dense matter EoS [see, e.g., 301, 265, 302, 303, for recent analyses that include PSR J04374715] and the magnetic field geometry of the MSPs [2].
@@ 2.1.2 MSP PPM with eXTP
NICER has a peak effective area at 1 keV of 1800 cm2. The eXTP SFA will be a factor of 2.5 larger below 2 keV (the energy range of interest for MSP PPM), with much lower background, enabling improved constraints on mass and radius for the NICER MSPs [304, 305, 306] and the collection of data sets large enough to enable PPM for MSPs that are too faint to be studied with NICER [307]. Figs. 3 and 4 show simulated pulse profiles and the level of constraints achievable within 1 Ms exposure times with eXTP for PSR J04374715 and PSR J0740+6620, compared to what is possible with current NICER data. Of greatest interest are nearby MSP binaries with precise measurements of the NS mass from radio pulse timing (see discussion in Sect. 6.1), since this reduces one source of uncertainty. In addition to the pulsars mentioned above already being studied by NICER, these include PSR J19093744 [ ; 308], PSR J07511807 [ ; 309], PSR J22220137 [ ; 310], and PSR J16142230 [ ; 308].
@@ 2.2.2 Combined constraints
The constraints on geometry derived from the polarimetric data for the AMXPs can then be used together with PPM on other (non-polarimetry) data sets. Fig. 7 shows simulated constraints for 100 ks of exposure time for eXTP-SFA for the scenario explored in Sect. 2.2.1, but using the geometric constraints derived from the eXTP-PFA analysis as priors. This analysis uses the AMXP PPM pipeline of [319], which incorporates an appropriate atmosphere model [318] and model for emission from the accretion disk. Compared to analysis of the same scenario for simulated NICER data, the constraints on mass and radius are much tighter with eXTP-SFA: the 68% credible interval on mass has reduced from 0.20 to 0.13 M⊙, and the 68% credible interval on radius has reduced from 0.19 to 0.11 km. The improvement is in part due to larger effective area, and in part due to polarization constraints. This is a conservative estimate for eXTP with an effective observation time here of only 100 ks, compared to the simulation in [319] of 130 ks.
@@ 3 Other-methods: spectroscopy, polarimetry and flux evolution
Thermonuclear X-ray bursts [e.g., 322, 323], also known as type I X-ray bursts, result from runaway thermonuclear burning on the surface of a NS in a LMXB system. The observational properties of type I X-ray bursts — such as their spectra, light curves, and peak luminosities — can be used to infer the NS mass and radius. This approach will be discussed in detail in Sect. 3.2.1.
During some type I X-ray bursts, X-ray spectra exhibit absorption edges at specific energy levels [e.g., 324, 325]. These features arise from the ionization of heavy elements in the NS atmosphere and provide valuable insights into the physical and chemical properties of the NS environment. This approach will be discussed in detail in Sect. 3.2.2.
Burst oscillations, hot spots generated by an as yet unknown mechanism during Type I X-ray bursts [see 326, for a review], may also provide information on the mass and radius. Pulse profile modeling for burst oscillations is challenging (given the intrinsic variability, the role of ongoing accretion, and uncertainty in the underlying surface pattern) [327, 328, 329]. However eXTP will observe burst oscillations with greater sensitivity in a softer waveband than existing observations, which may provide valuable insight into the mechanism and help to resolve surface pattern uncertainty.
@@ 3.4 Thermal evolution of NSs
While cooling processes tend to decrease the temperature of a NS, the relatively high temperatures observed in old NSs indicate the presence of significant heating mechanisms [e.g., 463]. Theoretical models must therefore incorporate not only cooling processes but also heating mechanisms, such as rotational energy dissipation (particularly important for MSPs), magnetic field decay (dominant in magnetars like SGR 1806-20), and vortex creep (relevant for older, slower pulsars).
@@ 4 Spin measurement
The sensitivity of a single 10-ks observation taken by eXTP-SFA can be calcualted in three cases: a faint source (0.5–10 keV flux of approximately ), a moderately bright one (), and a bright source (). These values correspond approximately to the fluxes observed in sources such as the AMXP XTE J1807–294 (faint), Aql X-1 (moderate), and Sco X-1 (bright). We adopted a coherent segment length of 512 seconds and targeted a signal-to-noise ratio of 5 for the pulsations. Under these conditions, we find that it is possible to place upper limits on the pulsed fraction of 0.9%, 0.3%, and 0.03% for faint, moderate, and bright sources, respectively.
@@ 6.1.2 Combined radio timing and optical (dynamical) observations
While radio timing provides the mass function and thereby constrains a combination of masses and orbital inclination, many NS binaries also contain an optically visible companion star whose dynamical signals can decisively break the degeneracy. In particular, the companion star can often be observed, supplying two key measurements, , and orbital inclination [583, and references therein]. Time‐resolved spectroscopy of the companion yields its radial‐velocity semi‐amplitude . One cross‐correlates absorption lines in each orbital phase with template stellar spectra, thereby measuring the Doppler shifts caused by the companion orbital motion [584, 585, 583]. Multi-band photometry of the companion over a full orbit also provides its brightness and temperature variations—crucial for determining the orbital inclination . In tight binaries such as “spider” pulsars (with a low-mass companion nearly filling its Roche lobe), the pulsar wind strongly heats one side of the companion, producing large brightness contrasts from day side to night side. Modeling these phase‐resolved fluxes (and accounting for effects like Roche‐lobe deformation, gravity darkening, and atmospheric limb darkening) tightly constrains [584, 585, 583].
@@ 6.2 Distance
For radio pulsars, a second type of parallax can be measured through timing observations, where distances are derived from variations in pulse arrival times caused by the curvature of the incoming wavefront as Earth moves in its orbit. Unlike imaging parallax, the precision of timing parallax is highest for pulsars at low ecliptic latitudes and lowest near the ecliptic pole [590, and references therein].
In addition, binary pulsar systems offer another method for distance determination through secular or annual variations in orbital parameters. However, these methods are typically limited to relatively nearby sources. Advances in telescope baseline lengths and sensitivity, such as those provided by the Square Kilometre Array (SKA), promise significant improvements in precision, enabling more accurate distance measurements [590, 591]. Another proposed method is the neutral hydrogren (H I) kinematic distance technique, uses the absorption of pulsar emission by Galactic H I gas, with distances inferred from a Galactic rotation model [592, and references therein]. Finally, the methods of distance measurements could be combined, such the timing, optical image and VLBI to improve the accuracy utilizing different facilities [590, 593].
@@ 7 X-ray study of dense matter in the multi-scale and multi-messenger era
The study of dense matter is a challenging task that requires theoretical models to extrapolate from matter at nuclear densities to the largely unknown regimes of high-density, neutron-rich or exotic compositions. Additionally, it depends on identifying the relevant degrees of freedom, which may range from nucleons to exotic particles, and even dark matter particles. One generally assumes that there is one theoretical model that can correctly explain the nuclear matter data of different physical situations obtained in both laboratory nuclear experiments and astronomical observations. It is necessary to combine efforts from different communities and discuss mutual interests and problems [594, 595]. Experimental nuclear physics provides critical data needed to benchmark theories of dense matter EOS residing in pulsars. The current and upcoming multimessenger observatories (e.g., LIGO/Virgo/KAGRA, FAST, SKA, LHAASO, HUBS, QTT) [596] will continue improving detection of NSs together with the precise measurements of their global properties, as well as their dynamical evolution. Laboratory experiments [597, 598, 599, 600, 601, 602] will provide an emerging understanding of dense matter EOS, together with the properties of strong interactions and the transition to deconfined quark matter. In this section, we outline recent progress and future perspectives along this line.
@@ 7.3 Interdisciplinary synthesis
Nevertheless, opposing results were obtained in a recent Bayesian analysis [625] that utilized experimental data on mid-rapidity proton observables — specifically, the elliptic flow () and the mean transverse kinetic energy () — in HICs with beam energies – GeV. Their sensitivity analysis for when all 15 available experimental data points are used inferred a broad peak structure in the squared speed of sound, indicating a rather stiff EOS up to four times the nuclear saturation density. However, if two key data points for at and GeV are excluded from the Bayesian inference — reducing the input to 13 data points — the extracted squared speed of sound drops drastically at high densities, consistent with a strong first-order phase transition.
The phase transition parameters derived from beta-stable NS matter can be tested in nearly symmetric nuclear matter produced in HIC [626]. However, to fully understand their implications, a deeper understanding of the connection between phase transitions in symmetric and asymmetric nuclear matter is required, and further investigations are needed on the roles played by temperature and magnetic field.
@@ 8 Summary
The synergy of eXTP instrumentation — combining large collecting area, polarimetry, and timing capabilities — allows for the deployment of multiple independent techniques. These include pulse profile modeling of MSPs, spectral modeling of thermonuclear bursts and their cooling tails, analysis of burst oscillations, and study of timing irregularities. In addition, eXTP will probe phenomena associated with nuclear burning on the NS surface (such as mHz QPOs) and accretion disk dynamics (such as kHz QPOs and relativistic Fe line profiles).
We anticipate that the comprehensive observational strategy of eXTP will open a new quantitative era for NS radius measurements. These measurements offer sensitivity to the nuclear symmetry energy, its density dependence, and possible phase transitions to exotic degrees of freedom (hyperons, quark matter) in the core. Additionally, a single SDD array in the SFA telescopes can be replaced by one pn-CCD, which offers combined spectral, timing, and imaging capabilities. The pn-CCD imaging performance (W50 =) enables tighter constraints on the phase-invariant component in energy-resolved pulse profiles through simultaneous observations of target sources. Furthermore, joint analysis of the SDD and pn-CCD data—for example in non-accreting MSPs—can enhance the precision of NS mass and radius measurements. Similar approaches have been employed in the analysis of NICER and XMM data (though not using simultaneous observations) using a joint log-likelihood function in pulse profile modeling.
@@ Acknowledgements.
AL is supported by the National Natural Science Foundation of China (grant No. 12273028). ALW, BD and TS acknowledge support from ERC Consolidator grant No. 865768 AEONS. ALW also acknowledges support from NWO grant ENW-XL OCENW.XL21.XL21.038. SG acknowledges the support of the CNES. S-NZ is supported by the National Natural Science Foundation of China (No. 12333007), the International Partnership Program of Chinese Academy of Sciences (No.113111KYSB20190020) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDA15020100). ZM is supported by the China national postdoctoral program for innovation talents (No. BX20240223) and the China Postdoctoral Science Foundation funded project (No. 2024M761948). YC acknowledges support from a Ramon y Cajal fellowship (RYC2021-032718-I) financed by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. XZ is supported by the Natural Science Foundation of Xinjiang Uygur Autonomous Region (No. 2023D01E20) National SKA Program of China (Grant No. 2020SKA0120300). XPZ and XZL are supported by the National Natural Science Foundation of China (grant No.12033001,12473039) and the WHW is supported by Zhejiang Provincial Natural Science Foundation of China under grant No. LQ24A030002. QC is supported by the National Natural Science Foundation of China (Grant No. 12003009). WWZ is supported by the National SKA Program of China No. 2020SKA0120200, the National Nature Science Foundation (grant No. 12041303). ZL is supported by the National Natural Science Foundation of China (grant No. 1227303). LS was supported by the National SKA Program of China (2020SKA0120300), the Beijing Natural Science Foundation (1242018), and the Max Planck Partner Group Program funded by the Max Planck Society. AD is supported by JSPS KAKENHI (the Japan Society for the Promotion of Science, Grants-in-Aid for Scientific Research) under grant No. 23K19056 and No. 25K17403. HS is supported by the National Natural Science Foundation of China (grant No. 12175109). CP acknowledges support from FCT (Fundação para a Ciência e a Tecnologia, I.P, Portugal) under projects UIDB/04564/2020 and 2022.06460.PTDC. LT is supported from the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, from the project PID2022-139427NB-I00 financed by the Spanish MCIN/AEI/10.13039/501100011033/FEDER, UE (FSE+), and by the CRC-TR 211 ’Strong-interaction matter under extreme conditions’- project Nr. 315477589 - TRR 211. LL is supported by the National Natural Science Foundation of China (grant No. 12122513). The work of KH, MM, AS and IS was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101020842) and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 279384907 – SFB 1245. ML acknowledges support from the ERC Consolidator grant No. 101002352 (LOVE-NEST). NR is supported by the European Research Council (ERC) via the Consolidator Grant “MAGNESIA” (No. 817661) and the Proof of Concept “DeepSpacePulse” (No. 101189496), by the Catalan grant SGR2021-01269, the Spanish grant ID2023-153099NA-I00, and by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M. AV is supported by the Research Council of Finland under grant No. 354533.

### v2

@@ 2 Pulse profile modeling
Pulse profile modeling (PPM) exploits the variation in brightness and spectrum as a function of the rotational phase to extract properties of either the emission or the NS itself. In the context of constraints on dense matter, PPM is used to model the emission of MSPs (rotation- or accretion- powered) [272]: more specifically the emission originating close to the NS surface, i.e., the emission affected the most by the strong NS surface gravity. As this is related to the NS compactness, PPM facilitates the measurement of the NS mass and radius, among other properties. This requires not only high quality pulse profile data (e.g., with order 10–100 sec timing resolution and 100 eV spectral resolution in the soft X-ray band) but also the ability to simulate pulse profiles using an adequate physical model (to form the basis for likelihood evaluations). Model ingredients include the physics of relativistic ray-tracing in the space-time of a rapidly rotating NS for a given mass and radius, a prescription for the possible properties of the X-ray emitting hotspots, an atmospheric beaming function, the effects of interstellar absorption and the instrument response (taking into account calibration uncertainty). Other relevant parameters will include the distance, pulsar inclination and any non-source background. Priors must be chosen for all of the model parameters. For a detailed overview of the simulation and inference process involved in PPM, see [273, 274] and references therein.
@@ 2.1.1 State of the art - MSP PPM with NICER
NICER is a soft X-ray telescope operating on the International Space Station since 2017. Its primary science goal is to explore the EOS of NSs by collecting high quality pulse profile data sets of MSPs to be used for PPM. The NICER MSPs are also radio pulsars: radio timing provides the spin ephemeris necessary to generate the X-ray pulse profile, and other informative priors such as distance or, if the source is in a binary, the mass and inclination [282, 283]. Constraints on background can come either from NICER background models (e.g., [284]) or indirectly via joint fitting of (phase-averaged) data for the same source from other X-ray observatories with imaging capabilities (e.g., XMM-Newton, with a well constrained background, enabling a good estimate of the number of photons coming from the source in the XMM-Newton data set).
To date, PPM using NICER data has yielded results for five MSPs: PSR J0030+0451, for which the mass and inclination are not known a priori [285, 286, 277, 287, 288, 289, 290]; the 2.1 pulsar PSR J0740+6620 [291, 292, 293, 294, 287, 295, 296, 297]; the 1.4 pulsar PSR J04374715 [281]; PSR J12311411, which has only a weakly-informative prior on mass and inclination [298, 275]; and the pulsar PSR J0614-3329 [299]. In support of this effort, extensive work has gone into developing and testing PPM pipelines [273, 274, 300, 301]. The tightest constraints on radius so far (for PSR J04374715, [281]) are at the % level (68% credible interval). The results are being used to place constraints on both the dense matter EoS [see, e.g., 302, 240, 303, 304, for recent analyses that include PSR J04374715] and the magnetic field geometry of the MSPs [2].
@@ 2.1.2 MSP PPM with eXTP
NICER has a peak effective area at 1 keV of 1800 cm2. The eXTP SFA will be a factor of 2.5 larger below 2 keV (the energy range of interest for MSP PPM), with much lower background, enabling improved constraints on mass and radius for the NICER MSPs [305, 306, 307] and the collection of data sets large enough to enable PPM for MSPs that are too faint to be studied with NICER [308]. Figs. 3 and 4 show simulated pulse profiles and the level of constraints achievable within 1 Ms exposure times with eXTP for PSR J04374715 and PSR J0740+6620, compared to what is possible with current NICER data. Of greatest interest are nearby MSP binaries with precise measurements of the NS mass from radio pulse timing (see discussion in Sect. 6.1), since this reduces one source of uncertainty. In addition to the pulsars mentioned above already being studied by NICER, these include PSR J19093744 [ ; 309], PSR J07511807 [ ; 310], PSR J22220137 [ ; 311], and PSR J16142230 [ ; 309].
@@ 2.2.1 Geometric constraints from polarimetry
| Observational Technique | Instrumental Mode | Required Effective Exposure | Expected Precision on M-R | Leading Systematic Limits |
| Pulse profile: non-accreting MSP | SFA long-term pointed observation | 1–2 Ms per source | Radius: 5–10%; Mass: 5–10% | Background, hot spot configuration, mass prior constraint |
| Pulse profile: accreting MSP | SFA and PFA long-term pointed observation | 100–600 ks per source | Radius: 5–10%; Mass: 10–15% | Hot spot and accretion disk modelling, accretion emission variability |
@@ 2.2.2 Combined constraints
The constraints on geometry derived from the polarimetric data for the AMXPs can then be used together with PPM on other (non-polarimetry) data sets. Fig. 7 shows simulated constraints for 100 ks of exposure time for eXTP-SFA for the scenario explored in Sect. 2.2.1, but using the geometric constraints derived from the eXTP-PFA analysis as priors. This analysis uses the AMXP PPM pipeline of [320], which incorporates an appropriate atmosphere model [318] and model for emission from the accretion disk. Compared to analysis of the same scenario for simulated NICER data, the constraints on mass and radius are much tighter with eXTP-SFA: the 68% credible interval on mass has reduced from 0.20 to 0.13 M⊙, and the 68% credible interval on radius has reduced from 0.19 to 0.11 km. The improvement is in part due to larger effective area, and in part due to polarization constraints. This is a conservative estimate for eXTP with an effective observation time here of only 100 ks, compared to the simulation in [320] of 130 ks. Table 1 summarizes the quantitative performance of PPM techniques in constraining NS mass and radius.
@@ 3 Other-methods: spectroscopy, polarimetry and flux evolution
Thermonuclear X-ray bursts [e.g., 323, 324], also known as type I X-ray bursts, result from runaway thermonuclear burning on the surface of a NS in a LMXB system. The observational properties of type I X-ray bursts — such as their spectra, light curves, and peak luminosities — can be used to infer the NS mass and radius. This approach will be discussed in detail in Sect. 3.2.1.
During some type I X-ray bursts, X-ray spectra exhibit absorption edges at specific energy levels [e.g., 325, 326]. These features arise from the ionization of heavy elements in the NS atmosphere and provide valuable insights into the physical and chemical properties of the NS environment. This approach will be discussed in detail in Sect. 3.2.2.
Burst oscillations, hot spots generated by an as yet unknown mechanism during Type I X-ray bursts [see 131, for a review], may also provide information on the mass and radius. PPM for burst oscillations is challenging (given the intrinsic variability, the role of ongoing accretion, and uncertainty in the underlying surface pattern) [327, 328, 329]. However eXTP will observe burst oscillations with greater sensitivity in a softer waveband than existing observations, which may provide valuable insight into the mechanism and help to resolve surface pattern uncertainty.
@@ 3.4 Thermal evolution of NSs
While cooling processes tend to decrease the temperature of a NS, the relatively high temperatures observed in old NSs indicate the presence of significant heating mechanisms [e.g., 461]. Theoretical models must therefore incorporate not only cooling processes but also heating mechanisms, such as rotational energy dissipation (particularly important for MSPs), magnetic field decay (dominant in magnetars like SGR 1806-20), and vortex creep (relevant for older, slower pulsars).
@@ 4 Spin measurement
The sensitivity of a single 10-ks observation taken by eXTP-SFA can be calculated in three cases: a faint source (0.5–10 keV flux of approximately ), a moderately bright one (), and a bright source (). These values correspond approximately to the fluxes observed in sources such as the AMXP XTE J1807–294 (faint), Aql X-1 (moderate), and Sco X-1 (bright). We adopted a coherent segment length of 512 seconds and targeted a signal-to-noise ratio of 5 for the pulsations. Under these conditions, we find that it is possible to place upper limits on the pulsed fraction of 0.9%, 0.3%, and 0.03% for faint, moderate, and bright sources, respectively.
@@ 6.1.2 Combined radio timing and optical (dynamical) observations
While radio timing provides the mass function and thereby constrains a combination of masses and orbital inclination, many NS binaries also contain an optically visible companion star whose dynamical signals can decisively break the degeneracy. In particular, the companion star can often be observed, supplying two key measurements, , and orbital inclination [575, and references therein]. Time‐resolved spectroscopy of the companion yields its radial‐velocity semi‐amplitude . One cross‐correlates absorption lines in each orbital phase with template stellar spectra, thereby measuring the Doppler shifts caused by the companion orbital motion [576, 577, 575]. Multi-band photometry of the companion over a full orbit also provides its brightness and temperature variations—crucial for determining the orbital inclination . In tight binaries such as “spider” pulsars (with a low-mass companion nearly filling its Roche lobe), the pulsar wind strongly heats one side of the companion, producing large brightness contrasts from day side to night side. Modeling these phase‐resolved fluxes (and accounting for effects like Roche‐lobe deformation, gravity darkening, and atmospheric limb darkening) tightly constrains [576, 577, 575].
@@ 6.2 Distance
For radio pulsars, a second type of parallax can be measured through timing observations, where distances are derived from variations in pulse arrival times caused by the curvature of the incoming wavefront as Earth moves in its orbit. Unlike imaging parallax, the precision of timing parallax is highest for pulsars at low ecliptic latitudes and lowest near the ecliptic pole [581, and references therein].
In addition, binary pulsar systems offer another method for distance determination through secular or annual variations in orbital parameters. However, these methods are typically limited to relatively nearby sources. Advances in telescope baseline lengths and sensitivity, such as those provided by the Square Kilometre Array (SKA), promise significant improvements in precision, enabling more accurate distance measurements [581, 582]. Another proposed method is the neutral hydrogren (H I) kinematic distance technique, uses the absorption of pulsar emission by Galactic H I gas, with distances inferred from a Galactic rotation model [583, and references therein]. Finally, the methods of distance measurements could be combined, such the timing, optical image and VLBI to improve the accuracy utilizing different facilities [581, 584].
@@ 7 X-ray study of dense matter in the multi-scale and multi-messenger era
The study of dense matter is a challenging task that requires theoretical models to extrapolate from matter at nuclear densities to the largely unknown regimes of high-density, neutron-rich or exotic compositions. Additionally, it depends on identifying the relevant degrees of freedom, which may range from nucleons to exotic particles, and even dark matter particles. One generally assumes that there is one theoretical model that can correctly explain the nuclear matter data of different physical situations obtained in both laboratory nuclear experiments and astronomical observations. It is necessary to combine efforts from different communities and discuss mutual interests and problems [lattimer2021, schatz2022horizons]. Experimental nuclear physics provides critical data needed to benchmark theories of dense matter EOS residing in pulsars. The current and upcoming multimessenger observatories (e.g., LIGO/Virgo/KAGRA, FAST, SKA, LHAASO, HUBS, QTT) [585] will continue improving detection of NSs together with the precise measurements of their global properties, as well as their dynamical evolution. Laboratory experiments [586, 587, 588, 589, 590, 591] will provide an emerging understanding of dense matter EOS, together with the properties of strong interactions and the transition to deconfined quark matter. In this section, we outline recent progress and future perspectives along this line.
@@ 7.3 Interdisciplinary synthesis
Nevertheless, opposing results were obtained in a recent Bayesian analysis [614] that utilized experimental data on mid-rapidity proton observables — specifically, the elliptic flow () and the mean transverse kinetic energy () — in HICs with beam energies – GeV. Their sensitivity analysis for when all 15 available experimental data points are used inferred a broad peak structure in the squared speed of sound, indicating a rather stiff EOS up to four times the nuclear saturation density. However, if two key data points for at and GeV are excluded from the Bayesian inference — reducing the input to 13 data points — the extracted squared speed of sound drops drastically at high densities, consistent with a strong first-order phase transition. The phase transition parameters derived from beta-stable NS matter can be tested in nearly symmetric nuclear matter produced in HIC [615]. However, to fully understand their implications, a deeper understanding of the connection between phase transitions in symmetric and asymmetric nuclear matter is required, and further investigations are needed on the roles played by temperature and magnetic field.
@@ 8 Summary
The synergy of eXTP instrumentation — combining large collecting area, polarimetry, and timing capabilities — allows for the deployment of multiple independent techniques. These include PPM of MSPs, spectral modeling of thermonuclear bursts and their cooling tails, analysis of burst oscillations, and study of timing irregularities. In addition, eXTP will probe phenomena associated with nuclear burning on the NS surface (such as mHz QPOs) and accretion disk dynamics (such as kHz QPOs and relativistic Fe line profiles).
We anticipate that the comprehensive observational strategy of eXTP will open a new quantitative era for NS radius measurements. These measurements offer sensitivity to the nuclear symmetry energy, its density dependence, and possible phase transitions to exotic degrees of freedom (hyperons, quark matter) in the core. Additionally, a single SDD array in the SFA telescopes can be replaced by one pn-CCD, which offers combined spectral, timing, and imaging capabilities. The pn-CCD imaging performance (W50 =) enables tighter constraints on the phase-invariant component in energy-resolved pulse profiles through simultaneous observations of target sources. Furthermore, joint analysis of the SDD and pn-CCD data—for example in non-accreting MSPs—can enhance the precision of NS mass and radius measurements. Similar approaches have been employed in the analysis of NICER and XMM data (though not using simultaneous observations) using a joint log-likelihood function in PPM.
@@
Acknowledgements. This work is supported by China’s Space Origins Exploration Program. AL is supported by the National Natural Science Foundation of China (grant No. 12273028). ALW, BD and TS acknowledge support from ERC Consolidator grant No. 865768 AEONS. ALW also acknowledges support from NWO grant ENW-XL OCENW.XL21.XL21.038. SG acknowledges the support of the CNES. S-NZ is supported by the National Natural Science Foundation of China (No. 12333007), the International Partnership Program of Chinese Academy of Sciences (No.113111KYSB20190020) and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDA15020100). ZM is supported by the China national postdoctoral program for innovation talents (No. BX20240223) and the China Postdoctoral Science Foundation funded project (No. 2024M761948). YC acknowledges support from a Ramon y Cajal fellowship (RYC2021-032718-I) financed by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. AP acknowledges support from grant PID2021-124581OB-I0, PID2024-155316NB-I00 and 2021SGR00426. XZ is supported by the Natural Science Foundation of Xinjiang Uygur Autonomous Region (No. 2023D01E20) National SKA Program of China (Grant No. 2020SKA0120300). XPZ and XZL are supported by the National Natural Science Foundation of China (grant No.12033001,12473039) and the WHW is supported by Zhejiang Provincial Natural Science Foundation of China under grant No. LQ24A030002. QC is supported by the National Natural Science Foundation of China (Grant No. 12003009). WWZ is supported by the National SKA Program of China No. 2020SKA0120200, the National Nature Science Foundation (grant No. 12041303). ZL is supported by the National Natural Science Foundation of China (grant No. 1227303). LS was supported by the National SKA Program of China (2020SKA0120300), the Beijing Natural Science Foundation (1242018), and the Max Planck Partner Group Program funded by the Max Planck Society. AD is supported by JSPS KAKENHI (the Japan Society for the Promotion of Science, Grants-in-Aid for Scientific Research) under grant No. 23K19056 and No. 25K17403. HS is supported by the National Natural Science Foundation of China (grant No. 12175109). CP acknowledges support from FCT (Fundação para a Ciência e a Tecnologia, I.P, Portugal) under projects UIDB/04564/2020 and 2022.06460.PTDC. LT is supported from the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, from the project PID2022-139427NB-I00 financed by the Spanish MCIN/AEI/10.13039/501100011033/FEDER, UE (FSE+), and by the CRC-TR 211 ’Strong-interaction matter under extreme conditions’- project Nr. 315477589 - TRR 211. LL is supported by the National Natural Science Foundation of China (grant No. 12122513). The work of KH, MM, AS and IS was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101020842) and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 279384907 – SFB 1245. ML acknowledges support from the ERC Consolidator grant No. 101002352 (LOVE-NEST). NR is supported by the European Research Council (ERC) via the Consolidator Grant “MAGNESIA” (No. 817661) and the Proof of Concept “DeepSpacePulse” (No. 101189496), by the Catalan grant SGR2021-01269, the Spanish grant ID2023-153099NA-I00, and by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M. AV is supported by the Research Council of Finland under grant No. 354533.
Conflict of interest. The authors declare that they have no conflict of interest.